# itemsep=1pt,topsep=2pt,parsep=1pt,partopsep=1pt

### v2

@@ 2.8. Notations
By writing $m=sN$ and $n=tN$ and by [Uch11, Corollary 1.1], the above quantity equals to
(3.7) $\begin{split}\sum\limits_{s<t\in\{\frac{1}{N},\cdots,\frac{N-1}{N}\}}\sum\limits_{u,v\in\frac{1}{\sqrt{N}}{\mathbbm{Z}}}&\frac{1}{\sqrt{N}}\varphi(u)\frac{1}{N}\big{(}Q(|u|,s)+R_{N}(|u|,s)\big{)}\\ &\times\frac{1+o(1)}{\sqrt{2\pi(t-s)}}\frac{1}{N}\Big{(}Q(|v|,1-t)+R_{N}(|v|,1-t)\Big{)}\frac{1}{\sqrt{N}}\psi(v).\end{split}$
The main term above, by a Riemann sum approximation, converges to
(3.8) $\mathscr{E}(\varphi,\psi)\!:=\!\frac{1}{(2\pi)^{\frac{3}{2}}}\!\iint_{0<s<t<1}\iint_{{\mathbbm{R}}\times{\mathbbm{R}}}\!\varphi(u)\frac{|uv|}{s\sqrt{s(t-s)(1-t)}(1-t)}e^{-\frac{1}{2}\left(\frac{u^{2}}{s}+\frac{v^{2}}{1-t}\right)}\psi(v){\mathrm{d}}u{\mathrm{d}}v{\mathrm{d}}t{\mathrm{d}}s.$
Note that by a change of variable $x=u/\sqrt{s}$ and $y=v/\sqrt{1-t}$, we have that
@@
where the finiteness can be found in [CSZ16, Lemma B.2]. Hence, $\mathscr{E}(\varphi,\psi)$ is well defined.
For the error term, we only treat the term with $u,s$. The term with $v,t$ is exactly the same. Note that if $|u|\leq\sqrt{s}$, then $(sN)^{\frac{3}{2}}\geq((u^{2}N)^{\frac{3}{2}}\vee 1)$ since $s\geq\frac{1}{N}$. The error term is then bounded by
(3.9) $\sum\limits_{u\in\frac{1}{\sqrt{N}}{\mathbbm{Z}}\atop|u|\leq\sqrt{s}}\frac{1}{\sqrt{N}}\varphi(u)|u|\frac{o(1)}{s^{\frac{3}{2}}N}+\sum\limits_{u\in\frac{1}{\sqrt{N}}{\mathbbm{Z}}\atop|u|>\sqrt{s}}\frac{1}{\sqrt{N}}\frac{\varphi(u)}{|u|^{2}}\frac{o(1)}{N}\leq\frac{o(1)}{s^{\frac{3}{2}}N}\int_{|u|\leq\sqrt{s}}|u|{\mathrm{d}}u+\frac{o(1)}{N}\int_{|u|\geq\sqrt{s}}\frac{1}{|u|^{2}}{\mathrm{d}}u=\frac{o(1)}{\sqrt{s}N},$
which is integrable with respect to $s$ at 00.
In summary, the error term is negligible, and it follows that
@@
Proof for Part (b). For $\mathscr{R}_{N}$ in (3.3), we show that $\lim_{N\to\infty}{\mathbbm{E}}[\mathscr{R}_{N}]=0$ and $\lim_{N\to\infty}{\rm Var}(\mathscr{R}_{N})=0$.
@@
(3.10) $\begin{split}{\mathbbm{E}}[\mathscr{R}_{N}]&=\sqrt{N}\int_{0}^{\frac{1}{N}}f(s){\mathrm{d}}s\sum\limits_{n=0}^{N-1}p_{n}(0)\int_{\frac{n}{N}}^{\frac{n+1}{N}}h(t){\mathrm{d}}t\sim\frac{1}{N}f(0)\int_{0}^{1}\frac{h(t)}{\sqrt{t}}{\mathrm{d}}t=\frac{O(1)}{N}.\end{split}$
For the variance, perform a polynomial chaos expansion to $Z_{0,n}^{\omega,\beta_{N}}$ (see Subsection 2.4) , and we note that ${\mathbbm{E}}[\mathscr{R}_{N}]$ is actually the deterministic term in $\mathscr{R}_{N}$. We have that
$\mathscr{R}_{N}-{\mathbbm{E}}[\mathscr{R}_{N}]=\sqrt{N}\int_{0}^{\frac{1}{N}}f(s){\mathrm{d}}s\sum\limits_{n=0}^{N-1}\sum\limits_{k=1}^{\infty}\sum\limits_{n_{0}:=0\leq n_{1}<\cdots<n_{k}=n}\zeta_{0}^{\mathbbm{1}_{\{n\neq 0\}}}\Big{(}\prod\limits_{j=1}^{k}p_{n_{j-1},n_{j}}(0)\zeta_{n_{j}}\Big{)}\int_{\frac{n}{N}}^{\frac{n+1}{N}}h(t){\mathrm{d}}t.$
@@
(3.11) $\begin{split}{\rm Var}(\mathscr{R}_{N})=N\Big{(}\int_{0}^{\frac{1}{N}}f(s){\mathrm{d}}s\Big{)}^{2}\sum\limits_{n=0}^{N-1}\sum\limits_{k=1}^{\infty}\big{(}\sigma_{N}^{2}\big{)}^{k+1}\sum\limits_{n_{0}:=0\leq n_{1}<\cdots<n_{k}=n}\Big{(}\prod_{j=1}^{k}u(n_{j}-n_{j-1})\Big{)}\Big{(}\int_{\frac{n}{N}}^{\frac{n+1}{N}}h(t){\mathrm{d}}t\Big{)}^{2},\end{split}$
@@
Recall the definition and properties of $\overline{U}_{N}(n)$ from (2.24)-(2.28). Then (3.11) is bounded above by
(3.12) $\begin{split}\frac{1}{N^{3}}\|f\|_{\infty}^{2}\|h\|_{\infty}^{2}\sum\limits_{n=0}^{N-1}\overline{U}(n)\leq\frac{C_{f,h}}{N^{3}}\sum\limits_{n=0}^{N-1}\frac{1}{N}G_{\vartheta}\Big{(}\frac{n}{N}\Big{)}\leq\frac{C}{N^{3}}\int_{0}^{1}G_{\vartheta}(t){\mathrm{d}}t=\frac{O(1)}{N^{3}}.\end{split}$
By (3.10) and (3.12), we have shown that $\mathscr{R}_{N}$ converges to 00 in $L_{2}$.
Proof for Part (c). We start by (i)$\Longrightarrow$(ii), which is more straightforward. Let
$Q_{N}(\varphi,s)$ $:=\int_{\mathbbm{R}}\varphi(x)Q_{N}(x,s){\mathrm{d}}x,$
$\widetilde{Q}_{N}(\psi,s)$ $:=\int_{\mathbbm{R}}\psi(-y)Q_{N}(-y,s){\mathrm{d}}y,$
$R_{N}(\varphi,s)$ $:=\int_{\mathbbm{R}}\varphi(x)R_{N}(x,s){\mathrm{d}}x,$
where $Q_{N}(x,s)$ and $R_{N}(x,s)$ have been introduced in (3.5). Then
(3.13) $\begin{split}&\sum\limits_{m=1}^{N-1}\sum\limits_{n=m}^{N-1}\Big{(}\int_{{\mathbbm{R}}}\varphi(x)\sqrt{N}q_{x_{N}}(m){\mathrm{d}}x\Big{)}\Big{(}\sqrt{N}Z^{\omega,\beta_{N}}_{m,n}\Big{)}\Big{(}\int_{{\mathbbm{R}}}\psi(x)\sqrt{N}q_{-y_{N}}(N-n){\mathrm{d}}y\Big{)}\\ =&\frac{1}{N^{2}}\sum_{s\leq t\in\frac{1}{N}{\mathbbm{Z}}\cap(0,1)}\Big{(}Q_{N}+R_{N}\Big{)}(\varphi,s)\Big{(}\sqrt{N}Z^{\omega,\beta_{N}}_{sN,tN}\Big{)}\Big{(}\widetilde{Q}_{N}+R_{N}\Big{)}(\psi,1-t).\\ \end{split}$
Assuming (i), the following weak convergence of random variables hold:
@@
which is indeed (1.27), where
$Q(\varphi,s):=\int_{\mathbbm{R}}\varphi(x)Q(x,s){\mathrm{d}}x\quad\text{and}\quad\widetilde{Q}(\psi,s):=\int_{\mathbbm{R}}\psi(-y)Q(-y,s){\mathrm{d}}y.$
Similar to (3.9), we can show that all error terms converge to 00. Hence, we conclude (i)$\Longrightarrow$(ii).
We then turn to (ii)$\Longrightarrow$(i). Note that by only knowing that $\widetilde{\mathcal{Z}}_{0,N}^{\beta_{N}}(\varphi,\psi)$ converges weakly, it is not straightforward to deduce (1.27). Hence, our strategy is to show that $Q(\varphi,s)$ or $\widetilde{Q}(\psi,s)$ are dense in $C([0,1])$, and then, to show ${\mathcal{Z}}_{N}^{\beta_{N}}(f,h)$ can be approximated by $\widetilde{\mathcal{Z}}_{0,N}^{\beta_{N}}(\varphi,\psi)-{\mathbbm{E}}[\widetilde{\mathcal{Z}}_{0,N}^{\beta_{N}}(\varphi,\psi)]$.
We only treat $Q(\varphi,s)$. The case $\widetilde{Q}(\psi,s)$ is exactly the same. Note that for $\varphi(x)=|x|^{2k+1}$,
(3.15) $\int_{{\mathbbm{R}}}\varphi(x)Q(x,s){\mathrm{d}}x=\int_{\mathbbm{R}}\frac{x^{2k+2}}{\sqrt{2\pi}s^{\frac{3}{2}}}e^{-\frac{x^{2}}{2s}}{\mathrm{d}}x=(2k+1)!!s^{k},$
which is a multiple of a power of $s$. By Weierstrass approximation theorem, for any $f\in C([0,1])$ and $\varepsilon>0$, we can choose $\widetilde{\varphi}_{\varepsilon}(x)$ of the form $\sum_{k=1}^{K_{\varepsilon}}C_{k}|x|^{2k+1}$ such that
$\sup_{s\in[0,1]}\Big{|}\int_{{\mathbbm{R}}}\widetilde{\varphi}_{\varepsilon}(x)Q(x,s){\mathrm{d}}x-f(x)\Big{|}<\varepsilon.$
Since $Q(x,s)$ has a double exponential decay in $x$, we can truncate and mollify $\widetilde{\varphi}_{\varepsilon}$ to obtain some $\varphi_{\varepsilon}\in C_{c}({\mathbbm{R}})$, such that
(3.16) $\sup\limits_{t\in[0,1]}\Big{|}\int_{{\mathbbm{R}}}\widetilde{\varphi}_{\varepsilon}(x)Q(x,s){\mathrm{d}}x-\int_{{\mathbbm{R}}}\varphi_{\varepsilon}(x)Q(x,s){\mathrm{d}}x\Big{|}<\varepsilon,$
and we have found the desired $\varphi_{\varepsilon}$.
To proceed, we write
(3.17) $\begin{split}{\mathcal{Z}}_{N}^{\beta_{N}}(f,h,\varphi,\psi):=&{\mathcal{Z}}_{N}^{\beta_{N}}(f,h)-\Big{(}\widetilde{\mathcal{Z}}_{0,N}^{\beta_{N}}(\varphi,\psi)-{\mathbbm{E}}\big{[}\widetilde{\mathcal{Z}}_{0,N}^{\beta_{N}}(\varphi,\psi)\big{]}\Big{)}.\end{split}$
We have the following estimate for (3.17), which concludes (ii)$\Longrightarrow$(i).
Lemma 3.1.
Given $\varepsilon>0$, for any $f,h\in C([0,1])$, we can find $\varphi,\psi\in C_{c}({\mathbbm{R}})$, such that
$\limsup\limits_{N\to\infty}{\mathbbm{E}}\Big{[}\big{(}{\mathcal{Z}}_{N}^{\beta_{N}}(f,h,\varphi,\psi)\big{)}^{2}\Big{]}\leq\varepsilon.$
Proof.
Write
$I_{m}^{(N)}(f):=I_{m}(f)=\int_{\frac{m}{N}}^{\frac{m+1}{N}}f(s){\mathrm{d}}s$
and we have that
@@
(3.18) $\begin{split}{\mathcal{Z}}_{N}^{\beta_{N}}(f,h,\varphi,\psi)=\sqrt{N}&\sum\limits_{m=1}^{N-1}\sum\limits_{n=m}^{N-1}I_{m}(f)\zeta_{m}^{\mathbbm{1}_{\{n\neq m\}}}\bigg{[}p_{m,n}(0)\zeta_{n}+\sum\limits_{k=1}^{\infty}\sum\limits_{n_{0}:=m\leq n_{1}<\cdots<n_{k}=n}\\ &\Big{(}\prod_{j=1}^{k}p_{n_{j-1},n_{j}}(0)\zeta_{n_{j}}\Big{)}\bigg{]}\bigg{(}I_{n}(h)-\widetilde{Q}_{N}\Big{(}\psi,\frac{N-n}{N}\Big{)}\bigg{)}\\ +\sqrt{N}&\sum\limits_{m=1}^{N-1}\sum\limits_{n=m}^{N-1}\bigg{(}I_{m}(f)-Q_{N}\Big{(}\varphi,\frac{m}{N}\Big{)}\bigg{)}\zeta_{m}^{\mathbbm{1}_{\{n\neq m\}}}\bigg{[}p_{m,n}(0)\zeta_{n}\\ &+\sum\limits_{k=1}^{\infty}\sum\limits_{n_{0}:=m\leq n_{1}<\cdots<n_{k}=n}\Big{(}\prod_{j=1}^{k}p_{n_{j-1},n_{j}}(0)\zeta_{n_{j}}\Big{)}I_{n}(h)\bigg{]}.\end{split}$
@@
Note that for any $\varepsilon>0$, we have shown that we can find $\varphi,\psi$, such that
(3.19) $\max\bigg{\{}~{}\bigg{|}I_{m}(f)-Q_{N}\Big{(}\varphi,\frac{m}{N}\Big{)}\bigg{|},\quad\bigg{|}I_{n}(h)-\widetilde{Q}_{N}\Big{(}\psi,\frac{N-n}{N}\Big{)}\bigg{|}~{}\bigg{\}}\leq\frac{\varepsilon}{N}$
uniformly in $11\leq m\leq n\leq N-$ 1, since $f,h$ have compact supports.
@@
Finally, by Proposition 2.3, the second moment of ${\mathcal{Z}}_{N}^{\beta_{N}}(f,h,\varphi,\psi)$ is bounded above by
@@
which concludes the proof. ∎
4. Moment estimates for ${\mathcal{Z}}_{0,N}^{\beta_{N}}(\varphi,\psi)$
In this subsection, we bound the higher central moments of $\widetilde{\mathcal{Z}}_{0,N}^{\beta_{N}}(\varphi,\psi)$. From now on, we work on $\widetilde{\mathcal{Z}}_{0,N}^{\beta_{N}}(\varphi,\psi)$ in (2.9). When there is no ambiguity, we omit the tilde in notations, and simply write ${\mathcal{Z}}_{0,N}^{\beta_{N}}(\varphi,\psi)$ by ${\mathcal{Z}}_{N}^{\beta_{N}}(\varphi,\psi)$ or ${\mathcal{Z}}_{N}$. We have the following Theorem 4.1. A technical novelty is that, when proving Proposition 4.3 below, we relax the assumption on the random walk to the vanishing third moment and a finite fourth moment, which requires a more careful analysis to utilize the local limit theorem.

### v3

@@ 2.8 Notations
By writing $m=sN$ and $n=tN$ and by [37, Corollary 1.1], the above quantity equals to
$\begin{split}\sum\limits_{s<t\in\{\frac{1}{N},\cdots,\frac{N-1}{N}\}}&\sum\limits_{u,v\in\frac{1}{\sqrt{N}}{\mathbbm{Z}}}\frac{1}{\sqrt{N}}\varphi(u)\frac{1}{N}\big(Q(|u|,s)+R_{N}(|u|,s)\big)\\ &\times\frac{1+o(1)}{\sqrt{2\pi(t-s)}}\frac{1}{N}\Big(Q(|v|,1-t)+R_{N}(|v|,1-t)\Big)\frac{1}{\sqrt{N}}\psi(v).\end{split}$ (3.8)
For the main term (involving $Q$ only), by a Riemann sum approximation, it converges to
$\begin{split}&\mathscr{E}(\varphi,\psi):=\\ \frac{1}{(2\pi)^{\frac{3}{2}}}\!\iint_{0<s<t<1}&\iint_{{\mathbbm{R}}\times{\mathbbm{R}}}\!\varphi(u)\frac{|uv|}{s\sqrt{s(t-s)(1-t)}(1-t)}e^{-\frac{1}{2}\left(\frac{u^{2}}{s}+\frac{v^{2}}{1-t}\right)}\psi(v){\mathrm{d}}u{\mathrm{d}}v{\mathrm{d}}t{\mathrm{d}}s.\end{split}$ (3.9)
Note that by a change of variables $x=u/\sqrt{s}$ and $y=v/\sqrt{1-t}$, we have that
@@
where the finiteness follows from [11, Lemma B.2]. Hence, $\mathscr{E}(\varphi,\psi)$ is well defined.
For the error terms (which involves $R_{N}$), we only treat the term with $u,s$. The other term with $v,t$ is exactly the same. Note that if $|u|\leq\sqrt{s}$, then $(sN)^{\frac{3}{2}}\geq((u^{2}N)^{\frac{3}{2}}\vee 1)$ since $s\geq\frac{1}{N}$. The error term is bounded by
$\begin{split}&\sum\limits_{u\in\frac{1}{\sqrt{N}}{\mathbbm{Z}}\atop|u|\leq\sqrt{s}}\frac{1}{\sqrt{N}}\varphi(u)|u|\frac{o(1)}{s^{\frac{3}{2}}N}+\sum\limits_{u\in\frac{1}{\sqrt{N}}{\mathbbm{Z}}\atop|u|>\sqrt{s}}\frac{1}{\sqrt{N}}\frac{\varphi(u)}{|u|^{2}}\frac{o(1)}{N}\\ \leq&\frac{o(1)}{s^{\frac{3}{2}}N}\int_{|u|\leq\sqrt{s}}|u|{\mathrm{d}}u+\frac{o(1)}{N}\int_{|u|\geq\sqrt{s}}\frac{1}{|u|^{2}}{\mathrm{d}}u=\frac{o(1)}{\sqrt{s}N},\end{split}$ (3.10)
which is integrable with respect to $s$ at 0. Hence, the error term is negligible.
By combining (3.9) and (3.10), it follows that
@@
Proof for Part (b). It suffices to show that for $\mathscr{R}_{N}$ in (3.3), $\lim_{N\to\infty}{\mathbbm{E}}[\mathscr{R}_{N}]=0$ and $\lim_{N\to\infty}{\rm Var}(\mathscr{R}_{N})=0$.
@@
$\begin{split}{\mathbbm{E}}[\mathscr{R}_{N}]&=\sqrt{N}\int_{0}^{\frac{1}{N}}f(s){\mathrm{d}}s\sum\limits_{n=0}^{N-1}p_{n}(0)\int_{\frac{n}{N}}^{\frac{n+1}{N}}h(t){\mathrm{d}}t\sim\frac{1}{N}f(0)\int_{0}^{1}\frac{h(t)}{\sqrt{t}}{\mathrm{d}}t=\frac{O(1)}{N}.\end{split}$ (3.11)
For the variance, perform a polynomial chaos expansion to $Z_{0,n}^{\omega,\beta_{N}}$ (see Subsection 2.4) , and note that ${\mathbbm{E}}[\mathscr{R}_{N}]$ is actually the deterministic term in the polynomial chaos expansion of $\mathscr{R}_{N}$. We have that
$\begin{split}&\mathscr{R}_{N}-{\mathbbm{E}}[\mathscr{R}_{N}]=\\ \sqrt{N}\int_{0}^{\frac{1}{N}}f(s){\mathrm{d}}s&\sum\limits_{n=0}^{N-1}\sum\limits_{k=1}^{\infty}\sum\limits_{n_{0}:=0\leq n_{1}<\cdots<n_{k}:=n}\zeta_{0}^{\mathbbm{1}_{\{n\neq 0\}}}\Big(\prod\limits_{j=1}^{k}p_{n_{j-1},n_{j}}(0)\zeta_{n_{j}}\Big)\int_{\frac{n}{N}}^{\frac{n+1}{N}}h(t){\mathrm{d}}t.\end{split}$
@@
$\begin{split}&{\rm Var}(\mathscr{R}_{N})=\\ N\Big(\int_{0}^{\frac{1}{N}}\!\!f(s){\mathrm{d}}s\Big)^{2}&\sum\limits_{n=0}^{N-1}\sum\limits_{k=1}^{\infty}\big(\sigma_{N}^{2}\big)^{k+1}\!\!\!\!\sum\limits_{n_{0}:=0\leq n_{1}<\cdots<n_{k}:=n}\!\!\Big(\prod_{j=1}^{k}u(n_{j}-n_{j-1})\Big)\!\Big(\int_{\frac{n}{N}}^{\frac{n+1}{N}}\!\!h(t){\mathrm{d}}t\Big)^{2},\end{split}$ (3.12)
@@
Recall the definition and properties of $\overline{U}_{N}(n)$ from (2.24)-(2.28). Then (3.12) is bounded from above by
$\begin{split}\frac{1}{N^{3}}\|f\|_{\infty}^{2}\|h\|_{\infty}^{2}\sum\limits_{n=0}^{N-1}\overline{U}(n)\leq\frac{C_{f,h}}{N^{3}}\sum\limits_{n=0}^{N-1}\frac{1}{N}G_{\vartheta}\Big(\frac{n}{N}\Big)\leq\frac{C}{N^{3}}\int_{0}^{1}G_{\vartheta}(t){\mathrm{d}}t=\frac{O(1)}{N^{3}}.\end{split}$ (3.13)
By Combining (3.11) and (3.13), we have shown that $\mathscr{R}_{N}$ converges to 0 in $L_{2}$.
Proof for Part (c). We only treat the case (3.4). The proof for the case (3.5) is totally the same. Recall (3.6) and note that we have proved in Part (a) that $R_{N}$ has a smaller order than $Q$ (see (3.10)). Hence, it suffice to treat
$\sum\limits_{u\in\frac{1}{\sqrt{N}}{\mathbbm{Z}}}\Big(\int_{u}^{u+\frac{1}{\sqrt{N}}}\varphi(x){\mathrm{d}}x\Big)\frac{1}{\sqrt{N}}Q\Big(u,\frac{m}{N}\Big)=\sum\limits_{u\in\frac{1}{\sqrt{N}}{\mathbbm{Z}}}\Big(\int_{u}^{u+\frac{1}{\sqrt{N}}}\varphi(x){\mathrm{d}}x\Big)\frac{|u|}{\sqrt{2\pi}m^{\frac{3}{2}}}e^{-\frac{(\sqrt{N}u)^{2}}{2m}}.$ (3.14)
Given $\varphi\in C_{c}({\mathbbm{R}})$, by its uniform continuity, (3.14) can be written by
$\frac{1}{\sqrt{N}}\sum\limits_{u\in\frac{1}{\sqrt{N}}{\mathbbm{Z}}}(\varphi(u)+o(1))\frac{|u|}{\sqrt{2\pi}m^{\frac{3}{2}}}e^{-\frac{(\sqrt{N}u)^{2}}{2m}}=\frac{1}{N}\sum\limits_{x\in{\mathbbm{Z}}}\Big(\varphi\Big(\frac{x}{\sqrt{N}}\Big)+o(1)\Big)\frac{|x|}{\sqrt{2\pi}m^{\frac{3}{2}}}e^{-\frac{x^{2}}{2m}}.$
Hence, by $\lim_{s\to 0}\frac{|x|}{s^{3/2}}\exp(-x^{2}/s)=0$ for all $x$, we can take
$f(s)=\sum\limits_{x\in{\mathbbm{Z}}}\varphi(0)\frac{|x|}{\sqrt{2\pi}s^{\frac{3}{2}}}e^{-\frac{x^{2}}{2s}}.$
Conversely, fix $f\in C([0,1])$. Note that for $\varphi(x)=|x|^{2k+1}$,
$\int_{{\mathbbm{R}}}\varphi(x)Q(x,s){\mathrm{d}}x=\int_{\mathbbm{R}}\frac{x^{2k+2}}{\sqrt{2\pi}s^{\frac{3}{2}}}e^{-\frac{x^{2}}{2s}}{\mathrm{d}}x=(2k+1)!!s^{k},$ (3.15)
which is a multiple of a power of $s$. By Weierstrass approximation theorem, for any $f\in C([0,1])$ and $\varepsilon>0$, we can choose $\widetilde{\varphi}_{\varepsilon}(x)$ of the form $\sum_{k=1}^{A_{\varepsilon}}C_{k}|x|^{2k+1}$ such that
$\sup_{s\in[0,1]}\Big|\int_{{\mathbbm{R}}}\widetilde{\varphi}_{\varepsilon}(x)Q(x,s){\mathrm{d}}x-f(x)\Big|<\varepsilon,$
where $A_{\varepsilon}$ is some positive integer. Since $Q(x,s)$ has a double exponential decay in $x$, we can truncate and mollify $\widetilde{\varphi}_{\varepsilon}$ to obtain some $\varphi_{\varepsilon}\in C_{c}({\mathbbm{R}})$, such that
$\sup\limits_{s\in[0,1]}\Big|\int_{{\mathbbm{R}}}\widetilde{\varphi}_{\varepsilon}(x)Q(x,s){\mathrm{d}}x-\int_{{\mathbbm{R}}}\varphi_{\varepsilon}(x)Q(x,s){\mathrm{d}}x\Big|<\varepsilon,$ (3.16)
and we have found the desired $\varphi_{\varepsilon}$.
@@
[…]
$Q(\varphi,s)$ $:=\int_{\mathbbm{R}}\varphi(x)Q(x,s){\mathrm{d}}x,$
$\widetilde{Q}(\psi,s)$ $:=\int_{\mathbbm{R}}\psi(-y)Q(-y,s){\mathrm{d}}y,$
$R_{N}(\varphi,s)$ $:=\int_{\mathbbm{R}}\varphi(x)R_{N}(x,s){\mathrm{d}}x.$
It is not hard to see (for example, in the proof for Part (c)) that $Q(\varphi,s)$ and $\widetilde{Q}(\psi,s)$ are continuous in $s$. To show (ii), namely the directed polymer partition function $\widetilde{{\mathcal{Z}}}_{0,N}^{\beta_{N}}(\varphi,\psi)$ in (3.1) converges, we note that the first term in the right-hand side converges due to Part (a). It remains to consider the second term
[…]
Assuming (i), the convergence of ${\mathcal{Z}}_{N}^{\beta_{N}}$ in (3.2) with $f(s)=Q(\varphi,s)$ and $h(t)=\widetilde{Q}(\psi,1-t)$ holds. Removing the term referring to $m=0$ thanks to Part (b), we have
[…]
@@
$\begin{array}[]{l}\displaystyle{\mathcal{Z}}_{N}^{\beta_{N}}(f,h,\varphi,\psi):=\sum\limits_{m=1}^{N-1}\sum\limits_{n=m}^{N-1}\Big(\int_{\frac{m}{N}}^{\frac{m+1}{N}}f(s){\mathrm{d}}s\Big)\Big(\sqrt{N}Z_{m,n}^{\omega,\beta_{N}}\Big)\Big(\int_{\frac{n}{N}}^{\frac{n+1}{N}}h(t){\mathrm{d}}t\Big)\\[5.0pt] \displaystyle\quad-\sum\limits_{m=1}^{N-1}\sum\limits_{n=m}^{N-1}\Big(\int_{{\mathbbm{R}}}\varphi(x)\sqrt{N}q_{x_{N}}(m){\mathrm{d}}x\Big)\Big(\sqrt{N}Z^{\omega,\beta_{N}}_{m,n}\Big)\Big(\int_{{\mathbbm{R}}}\psi(x)\sqrt{N}q_{-y_{N}}(N-n){\mathrm{d}}y\Big),\end{array}$ (3.19)
@@
We then show that given $\varepsilon>0$, for any $f,h\in C([0,1])$, we can find $\varphi,\psi\in C_{c}({\mathbbm{R}})$, such that
$\limsup\limits_{N\to\infty}{\mathbbm{E}}\Big[\big({\mathcal{Z}}_{N}^{\beta_{N}}(f,h,\varphi,\psi)\big)^{2}\Big]\leq\varepsilon.$ (3.20)
The weak convergence of ${\mathcal{Z}}_{N}^{\beta_{N}}(f,h)$ then follows from the weak convergence of $\widetilde{{\mathcal{Z}}}_{0,N}^{\beta_{N}}(\varphi,\psi)$, (3.20), and Part (a)-(b).
To prove (3.20), we write $I_{m}^{(N)}(f):=I_{m}(f)=\int_{\frac{m}{N}}^{\frac{m+1}{N}}f(s){\mathrm{d}}s$ and we have that
$\begin{split}&\quad\quad\quad{\mathcal{Z}}_{N}^{\beta_{N}}(f,h,\varphi,\psi)=\\ &\sqrt{N}\sum\limits_{m=1}^{N-1}\sum\limits_{n=m}^{N-1}I_{m}(f)\zeta_{m}^{\mathbbm{1}_{\{n\neq m\}}}\bigg[p_{m,n}(0)\zeta_{n}+\sum\limits_{k=1}^{\infty}\sum\limits_{n_{0}:=m\leq n_{1}<\cdots<n_{k}=n}\Big(\prod_{j=1}^{k}p_{n_{j-1},n_{j}}(0)\zeta_{n_{j}}\Big)\bigg]\\ &\times\bigg(I_{n}(h)-\widetilde{Q}_{N}\Big(\psi,\frac{N-n}{N}\Big)\bigg)+\sqrt{N}\sum\limits_{m=1}^{N-1}\sum\limits_{n=m}^{N-1}\bigg(I_{m}(f)-Q_{N}\Big(\varphi,\frac{m}{N}\Big)\bigg)\zeta_{m}^{\mathbbm{1}_{\{n\neq m\}}}\\ &\times\bigg[p_{m,n}(0)\zeta_{n}+\sum\limits_{k=1}^{\infty}\sum\limits_{n_{0}:=m\leq n_{1}<\cdots<n_{k}=n}\Big(\prod_{j=1}^{k}p_{n_{j-1},n_{j}}(0)\zeta_{n_{j}}\Big)I_{n}(h)\bigg].\end{split}$ (3.21)
For any $\varepsilon>0$ and large enough NN, we have shown that by (3.4) and (3.5), we can find $\varphi,\psi$, such that
$\max_{1\leq m\leq n\leq N-1}\bigg\{~\bigg|I_{m}(f)-Q_{N}\Big(\varphi,\frac{m}{N}\Big)\bigg|,\quad\bigg|I_{n}(h)-\widetilde{Q}_{N}\Big(\psi,\frac{N-n}{N}\Big)\bigg|~\bigg\}\leq\frac{\varepsilon}{N}.$ (3.22)
@@
Then by Proposition 2.3, the second moment of ${\mathcal{Z}}_{N}^{\beta_{N}}(f,h,\varphi,\psi)$ is bounded above by
@@
which concludes the proof.
4 Moment estimates for ${\mathcal{Z}}_{0,N}^{\beta_{N}}(\varphi,\psi)$
In this subsection, we bound the higher central moments of $\widetilde{\mathcal{Z}}_{0,N}^{\beta_{N}}(\varphi,\psi)$. From now on, we work on $\widetilde{\mathcal{Z}}_{0,N}^{\beta_{N}}(\varphi,\psi)$ in (2.9). When there is no ambiguity, we omit the tilde in notations, and simply write ${\mathcal{Z}}_{0,N}^{\beta_{N}}(\varphi,\psi)$ by ${\mathcal{Z}}_{N}^{\beta_{N}}(\varphi,\psi)$ or ${\mathcal{Z}}_{N}$. We have the following Theorem 4.1. A technical novelty is that, when proving Proposition 4.3 below, we relax the assumption on the random walk to (1.10), which requires a more careful analysis to utilize the local limit theorem.